OpenStudy (kmallette09):

Solve this quadratic equation using the quadratic formula. x 2 + 8x - 5 = 0

pooja195 (pooja195):

Hi there! Welcome To OpenStudy :-) Have you started the problem yet? If so what have you gotten to so far?

OpenStudy (will.h):

\[x^2 + 8x -5 =0\] we cannot use factoring by groups but we certainly can use the following rule \[-b+- \sqrt{b^2 -4ac} / 2a\] now we know that a= 1(leading coefficient) b=8 c=-5 therefore substitution method \[-8 +- \sqrt{8^2 -4(1)(-5)} /2(1)\] now we have 2 answers the 1st would be by the positive: \[-8+ \sqrt{8^2-4(1)(-5)}/2(1)\] Or the negative: \[-8-\sqrt{8^2-4(1)(-5)}/2(1)\] that's the answers.

OpenStudy (will.h):

make sense?

OpenStudy (kmallette09):

is that what x =

OpenStudy (will.h):

yes, there are 2 solutions. the 1st would be when positive comes after -b and the second would be negative comes after -b you may find the answers using calculator,

OpenStudy (will.h):

if you still confused i can explain more by details.

OpenStudy (kmallette09):

yes plz

OpenStudy (will.h):

okay. we have a rule that can be used to find the answers of s quadratic formula when we are unable to find the answers using the factoring by groups formula. that formula is called "the famous formula" which is. \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Then basically what you need to do is to substitue a=1 b=8 and c=-5

OpenStudy (kmallette09):

\[-8+5\sqrt{15} is that it

OpenStudy (will.h):

therefore. the 1st answer would be. \[\frac{ -b+\sqrt{b^2-4ac} }{ 2a}\] a=1 b=8 c=-5 \[\frac{ -8+ \sqrt{8^2-4(1)(-5)} }{ 2(1) }\] \[which equals \frac{ -8+\sqrt[2]{21} }{ 2 }\] which equals 0.58 that's the 1st answer tho

OpenStudy (kmallette09):

got it thanks

OpenStudy (will.h):

the second one would be in negative \[\frac{ -8-\sqrt{8^2-4(1)(-5)} }{ 2(1) }\] which equals -8.58

OpenStudy (will.h):

that's the answer. with details. :)

OpenStudy (mathmale):

Nice work, Will, very detailed. Next time, could you guide the other person through the problem solution so that he or she could find his or her own answers? That is the goal of OpenStudy (as you may know from having read the OpenStudy Code of Conduct).